OpenStudy (anonymous):

The sum of two consecutive integers is -46. What are the numbers?

OpenStudy (gorv):

let first number = x next number = x+1 x+x+1=-46

OpenStudy (gorv):

2x+1=-46

OpenStudy (gorv):

subtract 1 from both side

OpenStudy (gorv):

when u subtract u will found that sum of any two consecutive integer will never given u even sum it will alwys give you odd sum

OpenStudy (gorv):

like 2+3=5 odd 5+6=11 odd sum can not be even